OpenStudy (anonymous):

need help with this question 96oz_5lb which one is greater?

OpenStudy (anonymous):

If there are 16 oz in one pound, how many pounds is 96?

OpenStudy (anonymous):

6 pounds so the answer would be 96 oz.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

Your welcome

OpenStudy (anonymous):

i have 2 more questions.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

6 yard 1ft=

OpenStudy (anonymous):

6 yard x 1ft?

OpenStudy (anonymous):

i dont know how much is 6 yards i suck at math.

OpenStudy (anonymous):

there are three feet in one yard

OpenStudy (anonymous):

so 3*1 equals 3ft.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you so much last question.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

that last question you answer wasnt right the answer choices are 7 12 19 36

OpenStudy (anonymous):

19

OpenStudy (anonymous):

because 3*6=18+1=19

OpenStudy (anonymous):

ohh thats what i did wrong

OpenStudy (anonymous):

6/12 lb is_ oz

OpenStudy (anonymous):

12/6=2 16+16=_

OpenStudy (anonymous):

@purplechicken